OpenStudy (anonymous):

1. Simplify the following algebraic operation : 7x – 2(x-2) + 5(x+3)

OpenStudy (mew55):

7x-2x+4+5x+15=0 10x+19=0 10x=-19 x=-19/10

OpenStudy (anonymous):

thanks.

OpenStudy (mew55):

u r welcome